Question 10

Multigrain Health Foods Inc. is authorized to issue 5,000,000 common shares. In its initial public offering during 2010, Multigrain issued 500,000 common shares for $7.00 per share. Over the next year, Multigrain’s share price increased and the company issued 400,000 more shares at an average price of $8.50.

During the next seven years, from 2010 to 2016, Multigrain earned net income of $920,000. They declared and paid cash dividends of $140,000. A 10% stock dividend was distributed to the shareholders in 2016 on the shares outstanding. The market price was $8.00 per share when the stock dividend was distributed. At December 31, 2016, the company has total assets of $14,500,000 and total liabilities of $6,820,000.

  1.    Show the computation of Multigrain’s total shareholders’ equity at December 31, 2016.
  2.    Present a detailed computation of each element of shareholder’s equity.
Answer #1

Answer :

Computation of Number of shares, stock dividend, value of stock dividend

Number of shares = 500,000 + 400,000 = 900,000 Shares
Stock dividend = 10%*900,000 = 90,000 Shares
Value of stock dividned = 90000*8 = $720,000

Computation of Multigrain's total shareholders equity at December 31.2016

Particulars Amount ($)
Common stock capital -
500,000 shares @ $7 each 3,500,000.00
400,000 shares @ $8.5 each 3,400,000.00
Add : Stock dividend 90000 shares @ $8 each 720,000.00
A = 7,620,000.00
Multigrain Retained Earnings 2010 to 2016 Amount ($)
Net income 920,000.00
Less : Cash Dividend 140,000.00
= 780,000.00
Less : Stock dividend (90000 shares @ $8 each) 720,000.00
B = 60,000.00
Multigrain's Total shareholders equity at December 31.2016 = (A+B) = 7,680,000.00

Kindly Up-vote Thank You !!!